When a car is running with a velocity of 50 m/s, it is given a constant acceleration of 5 m/s². Find the displacement of the car 2 seconds later.

1
Expert's answer
2021-12-16T11:32:12-0500

"X=Vt+at^2/2=50\times 2+5\times 2^2/2=110m"